OpenStudy (anonymous):

medals rewarded! Use Cramer’s rule to solve the system of equations. x+4y-3z=9 -4x+8y-3z=36 4x+3y-3z=-3

OpenStudy (anonymous):

this takes work lots of work if you only want the answer, ask wolfram

OpenStudy (anonymous):

i will be happy to show you how to do it in wolfram if you like

OpenStudy (maddyx):

This is a pathetic one. Put this in a 3X3 determinant called A [1 4 -3] = [9] [-4 8 -3] = [36] [4 3 -3] = [-3] then, x = det (Ax) / det(A). det (Ax) is where you put the right hand determinant in the x column and so on. http://en.wikipedia.org/wiki/Cramer's_rule#Applications for more info

OpenStudy (anonymous):

ok, I'll try